### v1

@@ 4 Results
In situ $\gamma\alpha$ Toonen et al. (2018): $\alpha\lambda=2,\,\gamma=1.75$ Boissier & Prantzos (1999) $\sim$35 000 4 $1\times 10^{-4}$
In situ $\alpha\alpha$ Toonen et al. (2018): $\alpha\lambda=2$ Boissier & Prantzos (1999) $\sim$50 000 4 $8\times 10^{-5}$
In situ $\alpha\alpha 2$ Toonen et al. (2018): $\alpha\lambda=0.25$ Boissier & Prantzos (1999) $\sim$4 000 3 $8\times 10^{-4}$
@@
We present predictions for the number of MSP binaries in the Galactic bulge that could be detected by LISA, based on the population synthesis scenarios introduced in Section 3. Specifically, we explore two contrasting formation scenarios: one in which MSP binaries are deposited into the bulge through the accretion of disrupted globular clusters, and another in which they form in situ. The key assumptions and results of each model are summarised in Table 1. For each scenario, we report the distributions used to set binary properties, the total number of MSP binaries in the bulge ($N^{\rm bulge}_{\rm MSP}$), the number of binaries expected to be individually detectable by LISA ($N_{\rm MSP}^{\rm LISA}$), and the corresponding detection fraction ( $f^{\rm LISA}=N_{\rm MSP}^{\rm LISA}/N_{\rm MSP}^{\rm bulge}$). For the in situ models, the numbers of both the total bulge population and the detectable subset are estimated within the inner 1.7,kpc, chosen to enclose 95% of the bulge stars. For accreted models, detectability is assessed using Eq. (5), adopting a signal-to-noise ratio threshold of 7 for the nominal 4-year LISA mission duration. In this calculation, we also include the contribution of the Galactic gravitational-wave confusion foreground in the denominator, which is predominantly generated by the population of WDWD binaries. For the in situ models, the foreground has been estimated self-consistently in our previous work (Korol et al., 2024b), while for all accreted models we assume the foreground corresponding to the ‘in situ $\gamma\alpha$’ model, which is also adopted in the LISA Definition Study Report (Colpi et al., 2024).
@@
All models (except ‘accreted A’) yield at least one detectable system. The lack of detections in ‘accreted A’ arises from the smaller number of surviving systems at the present day, leaving no binaries above LISA’s detection threshold. The fraction of detectable MSP binaries lies in the range $10^{-5}$$10^{-4}$. Considering that LISA observations of Galactic binaries will be essentially complete at frequencies above a few millihertz (e.g., Korol et al., 2024a), even a single detection would indicate the presence of a much larger underlying population in the bulge—potentially numbering in the tens or hundreds of thousands. Conversely, a non-detection would begin to place upper limits on the size and properties of the bulge MSP binary population, helping to rule out the most optimistic models. Since both the accreted and in situ scenarios are likely to contribute to the real population, the fact that their predicted detection fractions are of similar order suggests that LISA could detect binaries from either origin. Even a handful of detections would not only confirm the presence of a significant MSP binary population in the bulge, but also provide valuable insights into its formation history.
@@
In Fig. 3, we present the detectable MSP binaries in the chirp mass–gravitational wave frequency parameter space overplotting all models listed in Table 1. It is evident that the two scenarios populate distinct regions of this space, particularly in chirp mass. This difference primarily reflects the assumed or resulting WD masses in each model. In the accreted scenario, the MSP is paired by construction with a He WD, whose mass is drawn from the Tauris & Savonije (1999) relation. As a result, the chirp masses cluster around $\sim$0.4 M. In contrast, in the in situ scenario, the companion mass emerges from the population synthesis. These binaries often contain more massive CO or ONe WD companions, resulting in higher chirp masses, typically in the range $\sim$0.7–1.1,M, because binaries with more massive companions are more likely to survive the common-envelope phase. This clearly separates the two formation channels in the chirp mass dimension and suggests that even a small number of LISA detections with well-measured chirp masses could help distinguish between different formation scenarios.
@@
We estimate the precision with which LISA can measure the parameters of MSPWD binaries using the Fisher matrix formalism (e.g., Cutler, 1998; Takahashi & Seto, 2002; Lau et al., 2020b). The gravitational wave frequency is the most robustly measured parameter, with a relative uncertainty given by:
$\frac{\sigma_{f_{\rm GW}}}{f_{\rm GW}}\simeq 8.7\times 10^{-7}\left(\frac{f_{\rm GW}}{2\,{\rm mHz}}\right)^{-1}\left(\frac{\rho}{10}\right)^{-1}\left(\frac{T_{\rm obs}}{4\,{\rm yr}}\right)^{-1}$ (13)
@@
The chirp mass, $\mathcal{M}$, can be inferred when both the frequency $f_{\rm GW}$ and its time derivative $\dot{f}_{\rm GW}$ are both measurable. These two quantities are related by:
@@
which allows one to invert for $\mathcal{M}$. The relative error on $\dot{f}_{\rm GW}$ is given by:
@@
and propagating uncertainties yields:
@@ 5 Discussion
In this study, we explored the prospects for detecting MSP binaries in the Galactic bulge with LISA. We considered two contrasting formation scenarios—accretion from disrupted globular clusters and in situ formation via standard binary evolution—and showed that both are expected to contribute individually resolvable sources. While these scenarios were chosen to bracket the range of plausible formation pathways for the bulge population, the true bulge population is likely to reflect a combination of both. Recent chemo-dynamical studies of the inner Milky Way suggest that a significant fraction of bulge stars formed early and rapidly, in burst-like episodes at high redshift (e.g. Belokurov & Kravtsov, 2022, 2023). Although our in situ models do not explicitly include a bursty star formation history, they are designed to capture early formation and may therefore already approximate key features of these more complex scenarios.
@@
Our accreted model, on the other hand, is intentionally data-driven and simple. It focuses exclusively on binaries consistent with having He WD companions, which constitute the majority ($\approx 70$% in the mass range $0.1$ – $0.5$ $M_{\mathrm{WD}}$) of the observed population (see Section 2). As a result, it likely underestimates the full detectable population from the accreted scenario. In particular, systems with more massive CO or ONe WD companions — though expected to be rarer — would exhibit higher chirp masses and thus be more readily detectable with LISA. Our detection estimates for the accreted scenario should therefore be interpreted as conservative lower limits. Future work could explore this in more detail using detailed $N$-body simulations of globular clusters.
@@
There is another MSP formation scenario that we have not considered in this study, in which bulge MSPs originate from the accretion-induced collapse of ONe WDs. In this channel, the WD collapses into a NS after accreting $\sim$0.1 M and surpassing the Chandrasekhar limit (e.g., Hurley et al., 2010; Tauris et al., 2013). This process is expected to produce negligible natal kicks, potentially enhancing NS retention in the bulge or within globular clusters. Recently, Gautam et al. (2022) proposed that the Galactic Center $\gamma$-ray excess can be naturally explained by a population of MSPs formed through this mechanism. However, MSPs formed via accretion-induced collapse are observationally difficult to distinguish from those produced via standard recycling, and theoretical uncertainties remain regarding the stability of mass accretion and the possibility of explosive outcomes before collapse.
@@ 5.1 Neutron star natal kicks
NSs are known to receive high velocities at birth due to asymmetric supernova explosions (Lyne & Lorimer, 1994). For a recent review see Popov et al. (2025). These natal kicks often reach hundreds of km s-1, well above the escape velocity of a typical globular cluster. Yet, the observed abundance of MSPs in globular clusters suggests that a non-negligible fraction of NSs are retained. Verbunt et al. (2017) and Igoshev (2020) analysed young radio pulsars with precise astrometric measurements and found that approximately $5\pm 3$% receive kicks low enough to be retained in globular clusters.
@@
Igoshev et al. (2021) further argued that analyses based solely on isolated pulsars are biased toward higher kick velocities, as NSs born with small kicks are more likely to remain in binaries and be detected as X-ray binaries or binary radio pulsars. This conclusion has been strengthened by recent studies of Be X-ray binaries, whose orbital eccentricities suggest a distinct low-kick component in the NS natal kick distribution (Valli et al., 2025). Specifically, this third component is consistent with a Maxwellian distribution with a standard deviation of just $\sim$5 km s-1—low enough that virtually all such NSs would be retained in globular clusters.
@@
In this study, we adopt a pragmatic approach: we assume that a sufficient number of NSs are retained in globular clusters, as supported by radio and $\gamma$-ray observations. Whether these NSs are retained due to receiving weak natal kicks or because they were born through accretion-induced collapse is not critical for our population synthesis. Both formation channels are consistent with our modelling as long as the resulting MSPWD binaries obey the Tauris & Savonije (1999) relation.
5.2 Ultra Compact X-ray Binaries
@@
UCXBs represent the evolutionary phase following detached NS/MSP-WD systems, once gravitational wave radiation shrinks the orbit sufficiently to trigger mass transfer (e.g., Nelemans & Jonker, 2010; Tauris, 2018). These systems can reach very short orbital periods (i.e., high gravitational wave frequencies) and can be detected by LISA in the bulge (see Section 2.2). For example, the highest-frequency UCXB in Fig. 2, 4U 1820-30, would be detectable by LISA in less than two years of observations, with its parameters well measured within the nominal four-year mission lifetime (Finch et al., 2023). While UCXBs typically experience accretion over longer timescales, their orbital periods increase over time, eventually moving out of LISA’s sensitivity band (see Fig. 2).
@@
We estimate the number of UCXBs in our accreted scenario by checking whether a given binary would merge within 2 Myr of its assigned age. We adopt this 2 Myr threshold based on detailed simulations by Tauris (2018), whose figure 2 shows that UCXBs spend approximately this duration with gravitational wave frequencies $f_{\mathrm{GW}}>3$ mHz. If a binary merges within 2 Myr, we assume the system is currently an UCXB. For binaries that do not meet this criterion, we compute their detectability by LISA as described in Section 2.2. We find that up to 3 and 4 UCXBs can be found in ‘Accreted C’ and ‘Accreted D’ models respectively if $N_{\mathrm{MSP}}^{\mathrm{bulge}}=1\times 10^{5}$. If $N_{\mathrm{MSP}}^{\mathrm{bulge}}=1.8\times 10^{4}$ we at most could find one UCXB.

### v2

@@ 3.1 Accreted scenario
In situ $\gamma\alpha$ Toonen et al. (2018): $\alpha\lambda=2,\,\gamma=1.75$ Boissier and Prantzos (1999) $\sim$35 000 4 $1\times 10^{-4}$
In situ $\alpha\alpha$ Toonen et al. (2018): $\alpha\lambda=2$ Boissier and Prantzos (1999) $\sim$50 000 4 $8\times 10^{-5}$
In situ $\alpha\alpha 2$ Toonen et al. (2018): $\alpha\lambda=0.25$ Boissier and Prantzos (1999) $\sim$4 000 3 $8\times 10^{-4}$
3.2 In situ scenario
In this study, we consider three models from Korol et al. (2024b) that adopt different prescriptions for the common-envelope phase—the phase of rapid orbital shrinkage that occurs when one star expands to fill its Roche lobe and engulfs its companion. Our model ‘in situ $\alpha\alpha$’ adopts the commonly used $\alpha$-formalism, which assumes that a fraction $\alpha$ of the orbital energy is used to eject the envelope, with the parameter $\lambda$ describing the envelope’s binding energy (Webbink, 1984). In this model, we assume a fixed value of $\alpha\lambda=2$. The model ‘in situ $\alpha\alpha 2$’ also uses the $\alpha$-formalism, but with a lower efficiency parameter, $\alpha\lambda=0.25$. In contrast, the model ‘in situ $\gamma\alpha$’ adopts an alternative prescription for the common-envelope phase, known as the $\gamma$-formalism (Nelemans et al., 2000), which is based on the conservation of angular momentum rather than energy. In this model, the $\gamma$-prescription is applied unless the binary contains a compact object or the common-envelope phase is triggered by a tidal instability (rather than dynamically unstable Roche-lobe overflow). As a result, the first common-envelope episode is typically treated using the $\gamma$-formalism, while the second—typically involving a giant donor and a WD or NS companion—is modelled using the $\alpha$-formalism (see Toonen et al., 2012). For this model, we adopt $\alpha\lambda=2$ and $\gamma=1.75$. Our choice of efficiency parameter values is guided by previous studies that calibrated them using observations. The $\alpha\alpha$ and $\gamma\alpha$ models follow the prescriptions constrained by observed double WD systems (hereafter WDWD; e.g. Nelemans et al., 2000; Nelemans and Tout, 2005; van der Sluys et al., 2006), while the lower efficiency in the $\alpha\alpha 2$ model is based on calibration from compact WDs in binaries with M-type main-sequence companions (Zorotovic et al., 2010; Toonen and Nelemans, 2013; Camacho et al., 2014; Zorotovic et al., 2014). In all three models, we adopt the NS natal kick prescription from Verbunt et al. (2017) as a distribution which describes observation of isolated radio pulsars well.
Our Milky Way model follows the approach of Nelemans et al. (2004), with the star formation rate linearly interpolated from the plane-projected grid computed by Boissier and Prantzos (1999), which is based on a spectro-photometric model of Galactic stellar populations. To represent the stellar bulge, we increase the star formation rate by a factor of two within the inner 3 kpc compared to the original Boissier and Prantzos (1999) prescription, and distribute the resulting binary population according to Eq. (12). This yields a present-day integrated stellar mass of approximately $2.6\times 10^{10}\,$M within this region.
4 Results
Figure 4: Detectable MSPWD binaries in the chirp mass–gravitational-wave frequency plane for both formation scenarios considered in this study. The in situ population is shown with orange stars, while systems from the accreted scenario are marked with green downward-pointing triangles. For comparison, the underlying population of detectable WDWD binaries is shown in blue-grey. Contours indicate regions of constant relative uncertainty in the chirp mass ($\sigma_{\mathcal{M}}/\mathcal{M}$) at levels of 1, 0.1, and 0.01, highlighting LISA’s ability to constrain binary parameters. The hatched region marks the part of parameter space where the chirp mass cannot be reliably measured ($\sigma_{\mathcal{M}}/\mathcal{M}>1$). The top axis shows the corresponding orbital period.
@@ 4 Results
We present predictions for the number of MSP binaries in the Galactic bulge that could be detected by LISA, based on the population synthesis scenarios introduced in Section 3. Specifically, we explore two contrasting formation scenarios: one in which MSP binaries are deposited into the bulge through the accretion of disrupted globular clusters, and another in which they form in situ. The key assumptions and results of each model are summarised in Table 3. For each scenario, we report the distributions used to set binary properties, the total number of MSP binaries in the bulge ($N^{\rm bulge}_{\rm MSP}$), the number of binaries expected to be individually detectable by LISA ($N_{\rm MSP}^{\rm LISA}$), and the corresponding detection fraction ( $f^{\rm LISA}=N_{\rm MSP}^{\rm LISA}/N_{\rm MSP}^{\rm bulge}$). For the in situ models, the numbers of both the total bulge population and the detectable subset are estimated within the inner 1.7 kpc, chosen to enclose 95% of the bulge stars. For accreted models, detectability is assessed using Eq. (5), adopting a signal-to-noise ratio threshold of 7 for the nominal 4-year LISA mission duration. In this calculation, we also include the contribution of the Galactic gravitational-wave confusion foreground in the denominator, which is predominantly generated by the population of WDWD binaries. For the in situ models, the foreground has been estimated self-consistently in our previous work (Korol et al., 2024b), while for all accreted models we assume the foreground corresponding to the ‘in situ $\gamma\alpha$’ model, which is also adopted in the LISA Definition Study Report (Colpi et al., 2024).
@@
All models (except ‘Accreted A’) yield at least one detectable system. The lack of detections in ‘Accreted A’ arises from the smaller number of surviving systems at the present day, leaving no binaries above LISA’s detection threshold. The fraction of detectable MSP binaries lies in the range $10^{-5}$$10^{-4}$. Considering that LISA observations of Galactic binaries will be essentially complete at frequencies above a few millihertz (e.g. Korol et al., 2024a), even a single detection would indicate the presence of a much larger underlying population in the bulge—potentially numbering in the tens or hundreds of thousands. Conversely, a non-detection would begin to place upper limits on the size and properties of the bulge MSP binary population, helping to rule out the most optimistic models. Since both the accreted and in situ scenarios are likely to contribute to the real population, the fact that their predicted detection fractions are of similar order suggests that LISA could detect binaries from either origin. Even a handful of detections would not only confirm the presence of a significant MSP binary population in the bulge, but also provide valuable insights into its formation history.
@@
In Fig. 4, we present the detectable MSP binaries in the chirp mass–-gravitational-wave frequency parameter space overplotting all models listed in Table 3. It is evident that the two scenarios populate distinct regions of this space, particularly in chirp mass. This difference primarily reflects the assumed or resulting WD masses in each model. In the accreted scenario, the MSP is paired by construction with a He WD, whose mass is drawn from the Tauris and Savonije (1999) relation. As a result, the chirp masses cluster around $\sim$0.4 M. In contrast, in the in situ scenario, the companion mass emerges from the population synthesis. These binaries often contain more massive CO or ONe WD companions, resulting in higher chirp masses, typically in the range $\sim$0.7–1.1 M, because binaries with more massive companions are more likely to survive the common-envelope phase. Although the separation between in the chirp mass produced in the two scenarios may appear very clear, we caution that it is partly driven by the simplified assumptions adopted in each case. Specifically, the accreted models enforce the Tauris and Savonije (1999) relation, producing only low-mass He WD companions, while the in situ models likely underproduce such extremely low-mass WDs owing to relatively weak magnetic braking prescriptions (e.g. Chen et al., 2021).
@@
We estimated the precision with which LISA can measure the parameters of MSPWD binaries using the Fisher matrix formalism (e.g. Cutler, 1998; Takahashi and Seto, 2002; Lau et al., 2020). The gravitational-wave frequency is the most robustly measured parameter, with a relative uncertainty given by
$\frac{\sigma_{f_{\rm GW}}}{f_{\rm GW}}\simeq 8.7\times 10^{-7}\left(\frac{f_{\rm GW}}{2\,{\rm mHz}}\right)^{-1}\left(\frac{\rho}{10}\right)^{-1}\left(\frac{T_{\rm obs}}{4\,{\rm yr}}\right)^{-1}.$ (13)
@@
The chirp mass, $\mathcal{M}$, can be inferred when both the frequency, $f_{\rm GW}$, and its time derivative, $\dot{f}_{\rm GW}$, are both measurable. These two quantities are related by
@@
which allows one to invert for $\mathcal{M}$. The relative error on $\dot{f}_{\rm GW}$ is given by
@@
, and propagating uncertainties yields
@@ 5 Discussion
In this study, we explored the prospects for detecting MSP binaries in the Galactic bulge with LISA. We considered two contrasting formation scenarios—accretion from disrupted globular clusters and in situ formation via standard binary evolution—and showed that both are expected to contribute individually resolvable sources. While these scenarios were chosen to bracket the range of plausible formation pathways for the bulge population, the true bulge population is likely to reflect a combination of both. Recent chemo-dynamical studies of the inner Milky Way suggest that a significant fraction of bulge stars formed early and rapidly, in burst-like episodes at high redshift (e.g. Belokurov and Kravtsov, 2022, 2023). Although our in situ models do not explicitly include a bursty star formation history, they are designed to capture early formation and may therefore already approximate key features of these more complex scenarios.
@@
Our accreted model is intentionally data-driven and simple. It focuses exclusively on binaries consistent with having He WD companions, which constitute the majority ($\approx 70$% in the mass range 0.10.10.50.5 $M_{\mathrm{WD}}$) of the observed MSP companions (see Section 2). As a result, it likely underestimates the full detectable population from the accreted scenario. In particular, systems with more massive CO or ONe WD companions — though expected to be rarer — would exhibit higher chirp masses and thus be more readily detectable with LISA. Our detection estimates for the accreted scenario should therefore be interpreted as conservative lower limits. Future work could explore this in more detail using detailed $N$-body simulations of globular clusters. On the other hand, extremely low-mass WDs are underrepresented in our in situ population synthesis because the models of Toonen et al. (2018), which underlie this scenario, adopt relatively weak magnetic braking prescriptions. As shown by Istrate et al. (2014) and Chen et al. (2021), stronger magnetic braking can produce more MSP binaries with extremely low-mass WD companion with short orbital periods. Consequently, the lack of detections around $\sim$0.4 M in the in situ scenario likely reflects this modelling limitation rather than an intrinsic absence of such systems.
@@
There is another MSP formation scenario that we have not considered in this study, in which bulge MSPs originate from the accretion-induced collapse of ONe WDs. In this channel, the WD collapses into a NS after accreting $\sim$0.1 M and surpassing the Chandrasekhar limit (e.g. Hurley et al., 2010; Tauris et al., 2013). This process is expected to produce negligible natal kicks, potentially enhancing NS retention in the bulge or within globular clusters. Recently, Gautam et al. (2022) proposed that the Galactic centre $\gamma$-ray excess can be naturally explained by a population of MSPs formed through this mechanism. However, MSPs formed via accretion-induced collapse are observationally difficult to distinguish from those produced via standard recycling, and theoretical uncertainties remain regarding the stability of mass accretion and the possibility of explosive outcomes before collapse.
5.1 Role of spider pulsars
Figure 1 shows that a noticeable fraction of $\gamma$-ray sources occupy the region associated with spider systems, that is, they belong either to the class of redbacks or black widows. These spider pulsars are typically identified through radio eclipses, which distinguish them from general MSPWD binaries.
In our population synthesis, we do not model whether the NS’s radio emission is eclipsed by its companion. In this respect, we represent redbacks as MSPs with He WD companions of masses below $0.4$ M, while black widows are not explicitly modelled because their formation pathways remain uncertain. For example, Chen et al. (2013) suggested that the fate of a binary depends primarily on the irradiation efficiency and beaming geometry of the NS, implying that redbacks do not evolve into black widows. On the other hand, Ginzburg and Quataert (2020) proposed that these two populations are evolutionarily linked.
As discussed above, LISA is expected to detect some WDNS binaries that may later be identified and characterised in the radio band as redbacks. While black widows could contribute to the $\gamma$-ray excess from the Galactic centre, they remain undetectable by LISA (see Fig. 3). Given that about 30 % of MSPs with detected $\gamma$-ray emission in the ATNF pulsar catalogue have median companion masses below 0.1 M (i.e. belonging to black widow category), our method may therefore underestimate the total number of WDNS-like systems in the Galactic bulge by up to $\sim$30 %.
@@ 5.2 Neutron star natal kicks
NSs are known to receive high velocities at birth due to asymmetric supernova explosions (Lyne and Lorimer, 1994). For a recent review see Popov et al. (2025). These natal kicks often reach hundreds of km s-1, well above the escape velocity of a typical globular cluster. Yet, the observed abundance of MSPs in globular clusters suggests that a non-negligible fraction of NSs are retained. Verbunt et al. (2017) and Igoshev (2020) analysed young radio pulsars with precise astrometric measurements and found that approximately $5\pm 3$% receive kicks low enough to be retained in globular clusters.
@@
Igoshev et al. (2021) further argued that analyses based solely on isolated pulsars are biased towards higher kick velocities, as NSs born with small kicks are more likely to remain in binaries and be detected as X-ray binaries or binary radio pulsars. This conclusion has been strengthened by recent studies of Be X-ray binaries, whose orbital eccentricities suggest a distinct low-kick component in the NS natal kick distribution (Valli et al., 2025). Specifically, this third component is consistent with a Maxwellian distribution with a standard deviation of just $\sim$5 km s-1—low enough that virtually all such NSs would be retained in globular clusters.
@@
In this study, we adopted a pragmatic approach: we assumed that a sufficient number of NSs are retained in globular clusters, as supported by radio and $\gamma$-ray observations. Whether these NSs are retained due to receiving weak natal kicks or because they were born through accretion-induced collapse is not critical for our population synthesis. Both formation channels are consistent with our modelling as long as the resulting MSPWD binaries obey the Tauris and Savonije (1999) relation.
5.3 Ultra-compact X-ray binaries
@@
The UCXBs represent the evolutionary phase following detached NSs (or MSPs) in binary systems with WD companions, once gravitational-wave radiation shrinks the orbit sufficiently to trigger mass transfer (e.g. Nelemans and Jonker, 2010; Tauris, 2018). These systems can reach very short orbital periods (i.e. high gravitational-wave frequencies) and can be detected by LISA in the bulge (see Section 2.2). For example, the highest-frequency UCXB in Fig. 3, 4U 1820-30, would be detectable by LISA in less than two years of observations, with its parameters well measured within the nominal four-year mission lifetime (Finch et al., 2023). While UCXBs typically experience accretion over longer timescales, their orbital periods increase over time, eventually moving out of LISA’s sensitivity band (see Fig. 3).
@@
We estimate the number of UCXBs in our accreted scenario by checking whether a given binary would merge within 2 Myr of its assigned age. We adopt this 2 Myr threshold based on detailed simulations by Tauris (2018), whose figure 2 shows that UCXBs spend approximately this duration with gravitational-wave frequencies $f_{\mathrm{GW}}>3$ mHz. If a binary merges within 2 Myr, we assume the system is currently an UCXB. For binaries that do not meet this criterion, we compute their detectability by LISA as described in Section 2.2. We find that up to three and four UCXBs can be found in ‘Accreted C’ and ‘Accreted D’ models, respectively, if $N_{\mathrm{MSP}}^{\mathrm{bulge}}=1\times 10^{5}$. For $N_{\rm MSP}^{\rm bulge}=1.8\times 10^{4}$, at most one UCXB is expected.